please tell me the way on how to solve this step-by-step by finding the roots of this in the form of quadratic equation.
A=160 meter square
L=2x+3
W=x

Answer by rothauserc(4718)

L * W = A where L is length, W is width, A is the area
(2x+3) * (x) = 160
2x^2 + 3x = 160
2x^2 + 3x - 160 = 0
*******************************************************
use quadratic formula
(-b + or - square root(b^2 - 4ac)) / 2a
x = (-3 + square root(9 - 4*2*(-160))) / 4 = 320
x = (-3 - square root(9 - 4*2*(-160))) / 4 = −321.5
*******************************************************
there are two solutions
x = 320 and x = -321.5